HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025  Not logged in ELOG logo
Message ID: 102     Entry time: Thu Apr 3 01:19:44 2025
Author: Zac, Johan 
Category: General 
Subject: Beam fluctuations 
Beam gone again between 01:14 and 01:18, then again at 01:19. HKR says there are problems with the source.
ELOG V3.1.5-fc6679b